# A Bayesian Estimator of Sample Size

### v2

@@ 7 Discussion
In this work we propose BESS as a new and simple Bayesian sample size estimator. BESS leads to a straightforward interpretation of estimated sample size: if the observed data exhibits certain level of evidence $e$ that supports the alternative hypothesis, with sample size $n$ one can conclude the alternative with confidence $c$, measured by posterior probability of the alternative. The statement is coherent with a subsequent Bayesian analysis of the trial using $n$ as the sample size. If the observed trial data exhibits evidence $e$, the posterior probability of alternative will be greater than $c$, corroborating with the sample size estimation. For dose optimization trials, the BESS SSR Cap design shows a superior performance under both Frequentist and Bayesian properties.
Simulation results show that for matched type I/II error rates and using vague priors, BESS produces similar sample size estimates as the standard Frequentist sample size estimation, even though the two are based on different philosophies and metrics. In addition, the preliminary comparison of various designs based on BESS suggests a slight advantage of the Bayesian approach.
@@ A.1 Posterior Probability of $H_{1}$ for One-arm Trial
Assume $\theta_{0}$ is a known constant, from Table (1) and equation (5), we have the posterior $f(\theta_{1}|\bm{y}_{n})$ in equation (3) for the three outcomes as follows:
@@
    $f(\theta_{1}|\bm{y}_{n})=\frac{\theta_{1}^{a+n\bar{y}}(1-\theta_{1})^{b+n(1-\bar{y})}}{B(a+n\bar{y},b+n(1-\bar{y}))}=\frac{\theta_{1}^{a+n[e+\theta_{0}]}(1-\theta_{1})^{b+n(1-[e+\theta_{0}])}}{B(a+n[e+\theta_{0}],b+n(1-[e+\theta_{0}]))}=f(\theta_{1}|e,n),$
@@ A.2 Posterior Probability of $H_{1}$ for Two-arm Trial with Continuous outcome and known variance
$f(\theta|\bm{y}_{n}^{d})=\frac{1}{\sqrt{\frac{2\pi}{\frac{1}{b}+\frac{n}{\sigma^{2}}}}}\exp{\left\{-\frac{1}{2}\left[\frac{1}{b}+\frac{n}{\sigma^{2}}\right]\left(\theta_{1}-\frac{1}{\frac{1}{b}+\frac{n}{\sigma^{2}}}\left[\frac{a}{b}+\frac{n\bar{y}^{d}}{\sigma^{2}}\right]\right)^{2}\right\}}$
$=\frac{1}{\sqrt{\frac{2\pi}{\frac{1}{b}+\frac{n}{\sigma^{2}}}}}\exp{\left\{-\frac{1}{2}\left[\frac{1}{b}+\frac{n}{\sigma^{2}}\right]\left(\theta_{1}-\frac{1}{\frac{1}{b}+\frac{n}{\sigma^{2}}}\left[\frac{a}{b}+\frac{ne}{\sigma^{2}}\right]\right)^{2}\right\}}=f(\theta_{1}|e,n),$
@@ A.4 BESS Algorithm 2’
A.5 Simulation Parameters for Coherence in Section 4.2
@@ A.6 Additional Simulation Setup and Results in Section 5
A.8 BESS Software and Verification
Package description
@@ A.8 BESS Software and Verification
    • Description: Compute the posterior probability of alternative (equation (3)) for the three outcome types of one arm trial.
    • Output: Posterior probability of $H_{1}$ (confidence) given sample size $n$, evidence $e$, and clinically minimum effect size $\theta^{*}$.
@@
    • Description: BESS algorithm 2 for binary outcomes in two-arm trial.
    • Output: Sample size per arm for binary outcomes in two-arm trial.
@@
    • Description: BESS algorithm 1 for continuous outcomes with known variance in two-arm trial.
    • Output: Sample size per arm for continuous outcomes with known variance in two-arm trial.
@@
    • Description: Compute the posterior probability of alternative (equation (3)) for binary outcome in two-arm trial.
    • Output: Posterior probability of $H_{1}$ (confidence) given sample size $n$, responders in treatment and control arms, and clinically minimum effect size $\theta^{*}$.
@@
    • Description: Compute the posterior probability of alternative (equation (3)) for continuous outcome in two-arm trial.
    • Output: Posterior probability of $H_{1}$ (confidence) given sample size $n$, difference in mean responses between treatment and control arms, and clinically minimum effect size $\theta^{*}$.
@@
    • Description: Compute the posterior probability of alternative (equation (3)) for count outcome in two-arm trial.
    • Output: Posterior probability of $H_{1}$ (confidence) given sample size $n$, event rates in treatment and control arms, and clinically minimum effect size $\theta^{*}$.
@@
Package validation
The results in the main manuscript verified the coherence property for two-arm trials. Moreover, an example which compares BESS sample size to the Standard SSE based on matched Type I/II error rates is demonstrated for binary outcome in a two-arm trial. The four designs functions are also showed in the main manuscript in Section 6.2. Here, we verify the coherence property and demonstrate similar sample sizes between BESS and standard SSE for binary outcome in one-arm trial as an example for one-arm trials.
To show the coherence property of one-arm trials, we first use BESS to estimate the sample sizes for the three outcomes in one-arm trial, with $e=0.4$, $\theta_{0}=0$, $\theta^{*}=0.3$, and $c=0.8$. We obtained 15 patients for binary outcome, 71 for continuous outcome, and 25 for count outcome. Next, following Section 4.2, we obtain a similar plot as Figure 2 in the manuscript, shown in Figure A.2, for the three outcomes in one-arm trial. The result suggests the coherence property of BESS for one-arm trial, which verifies the one-arm trial functions.
The following code segments uses BESS to estimate the sample sizes and produces Figure A.2:
set.seed(1234)
theta1 <- 0.4
theta0 <- 0
theta_s <- 0.3
e <- 0.4
c <- 0.8
> BESS_one_arm(theta_s, c, e, 1, 0.5, 0.5)\\$n
[1] 15
> BESS_one_arm(theta_s, c, e, 2, 0, 10)\\$n
[1] 71
> BESS_one_arm(theta_s, c, e, 3, 1, 2)\\$n
[1] 25
show_cohorence <- function(theta1, theta0, theta_s, e, c, n, a, b,
outcome, subv, main, sig = 1, q = 0.5,
n_sim = 10000){
sim_arr <- matrix(rep(NA, 2*n_sim),ncol=2)
for(i in 1:n_sim){
if(outcome == 1){
yi_trt <- rbinom(n, 1, prob = theta1)
yd <- mean(yi_trt) - theta0
}else if(outcome == 2){
yi_trt <- rnorm(n, mean = theta1, sd = sig)
yd <- mean(yi_trt) - theta0
}else{
yi_trt <- rpois(n, lambda = theta1)
yd <- mean(yi_trt) - theta0
post_H1 <- post_prob_H1_1a(n, theta_s, yd, outcome,
a, b, sig = sig, q = q)
sim_arr[i,] <- c(yd,post_H1)
smoothScatter(sim_arr[,1], sim_arr[,2],
[…]
ylab = ”Confidence”, bandwidth = 0.007,
nrpoints = 0, main = main)
abline(v = e, col = ”black”, lty = 2)
abline(h = c, col = ”red”, lty = 3)
show_cohorence(theta1, theta0, theta_s, e, c, 15, 0.5, 0.5, 1,
”(a)”, ”Binary outcome”)
show_cohorence(theta1, theta0, theta_s, e, c, 71, 0, 10, 2,
”(b)”, ”Continuous outcome”)
show_cohorence(theta1, theta0, theta_s, e, c, 25, 1, 2, 3,
”(c)”, ”Count-data outcome”)
Lastly, we verify the BESS methods for each trial and outcome types and show that similar sample sizes may be obtained with BESS and with standard SSE when the Type I and Type II error rates are matched. Following the simulation process in Section 5, the following code segments computes the Type I error rate and power (1 - Type II error rate) through simulation for one arm trials:
c, n, a, b, outcome = 1, sig = 1, q = 0.5,
n_sim = 10000){
result_arr <- matrix(rep(NA, 2*n_sim),ncol=2)
for(i in 1:n_sim){
if(outcome == 1){
yi_trtH1 <- rbinom(n, 1, prob = theta1H1)
yi_trtH0 <- rbinom(n, 1, prob = theta1H0)
}else if(outcome == 2){
yi_trtH1 <- rnorm(n, mean = theta1H1, sd = sig)
yi_trtH0 <- rnorm(n, mean = theta1H0, sd = sig)
}else{
yi_trtH1 <- rpois(n, lambda = theta1H1)
yi_trtH0 <- rpois(n, lambda = theta1H0)
ydH1 <- mean(yi_trtH1) - theta0
ydH0 <- mean(yi_trtH0) - theta0
post_H1_H1 <- post_prob_H1_1a(n, theta_s, ydH1, outcome,
a, b, sig = sig, q = q)
dec_H1 <- 0
if(post_H1_H1 >= c){
dec_H1 <- 1
post_H1_H0 <- post_prob_H1_1a(n, theta_s, ydH0, outcome,
a, b, sig = sig, q = q)
dec_H0 <- 0
if(post_H1_H0 >= c){
dec_H0 <- 1
result_arr[i,] <- c(dec_H0, dec_H1)
TIER <- sum(result_arr[,1] == 1)/n_sim
power <- sum(result_arr[,2] == 1)/n_sim
return(c(TIER, power))
> freq_prop_one_arm(0.4, 0.3, 0, 0.3, 0.8, 60,
0.5, 0.5, outcome = 1, n_sim = 1000)
[1] 0.2363 0.8278
For two-arm trials, the control arm’s data is also drawn from their respective distributions. The Frequentist sample size is then estimated using standard sample size packages such as “TrialSize” or online software. Table A.3 shows the estimated sample sizes via BESS and the oracle sample size with the standard SSE method. Again, the observation is similar to Table 2 in the manuscript, and further verifies these functions.
Trail Outcome Evidence Confidence type I power BESS Standard SSE
type type $e$ $c$ error rate $\alpha$ $1-\beta$ $n$ $n$
One-arm Binary 0.35 0.8 0.24 0.83 60 63
Continuous 0.4 0.8 0.20 0.50 71 71
Count 0.375 0.8 0.20 0.64 40 41
Two-arm Binary 0.20 0.9 0.11 0.38 35 36
Continuous 0.15 0.8 0.20 0.40 32 33
Count 0.3 0.7 0.27 0.88 51 51

### v3

@@ 7 Discussion
In this work we propose BESS as a new and simple Bayesian sample size estimator. BESS leads to a straightforward interpretation of estimated sample size: if the observed data exhibits certain level of evidence $e$ that supports the alternative hypothesis, with sample size $n$ one can conclude the alternative with confidence $c$, measured by posterior probability of the alternative. The statement is coherent with a subsequent Bayesian analysis of the trial using $n$ as the sample size. If the observed trial data exhibits evidence $e$, the posterior probability of alternative will be greater than $c$, corroborating with the sample size estimation. For dose optimization trials, the BESS SSR Cap design shows a superior performance under both Frequentist and Bayesian properties. Furthermore, for continuous outcomes, this design enables sample size re-estimation at interim analysis, allowing BESS to leverage the observed sample variance for more precise sample size estimation.
Simulation results show that for matched type I/II error rates and using vague priors, BESS produces similar sample size estimates as the standard Frequentist sample size estimation, even though the two are based on different philosophies and metrics. In addition, the preliminary comparison of various designs based on BESS suggests a slight advantage of the Bayesian approach, requiring fewer patients to achieve comparable error and false positive/negative rates.
In this work, we have implemented BESS for three outcome types: binary, continuous (assuming known variance), and count-data, for superiority and non-inferiority tests. However, as a framework, BESS is flexible enough to accommodate different outcome types and hypothesis. For example, BESS can be extended to composite outcomes. Suppose $y_{1}$ and $y_{2}$ represent two outcome types (categorical and continuous) under a sampling model $f(y_{1}|\theta_{1})f(y_{2}|\theta_{2})$ with a prior $\pi(\theta_{1},\theta_{2})$. In this case, BESS can be extended based on the confidence level of $\text{Pr}((\theta_{1},\theta_{2})\in\Theta^{*}|data)$, where $\Theta^{*}$ is a promising zone that reflects the desirable treatment effects defined based on both $\theta$’s. Additionally, BESS can be adapted to other hypothesis tests, such as equivalence testing. For instance, when the parameter is on the boundary under the null, the hypothesis is $H_{0}:\theta=\theta^{*}\text{ vs. }H_{1}:\theta\neq\theta^{*}$. BESS remains applicable in this setting but requires different priors. For example, under $H_{0}$, a point null can be used for $\theta_{1}$, such that $\theta_{0}|a_{0},b_{0},H=H_{0}\sim\frac{1}{C_{0}}Beta(\theta_{0}|a_{0},b_{0})I(\theta_{0}<1-\theta^{*})$, and $f(\theta_{1}|\theta_{0},H=H_{0})=I(\theta_{1}=\theta_{0}+\theta^{*})$). This modification leads to different formula of Equation (3).
@@ A.1 Posterior Probability of $H_{1}$ for One-arm Trial
Assume $\theta_{0}$ is a known constant, the density function $f(\bm{\theta}|\bm{y}_{n})$ in Equation (3) reduces to the posterior $f(\theta_{1}|\bm{y}_{n})$. From Table (1) and equation (5), we have the posterior $f(\theta_{1}|\bm{y}_{n})$ in equation (3) for the three outcomes as follows:
@@
    $f(\theta_{1}|\bm{y}_{n})=\frac{\theta_{1}^{a+n\bar{y}-1}(1-\theta_{1})^{b+n(1-\bar{y})-1}}{B(a+n\bar{y},b+n(1-\bar{y}))}=\frac{\theta_{1}^{a+n[e+\theta_{0}]-1}(1-\theta_{1})^{b+n(1-[e+\theta_{0}])-1}}{B(a+n[e+\theta_{0}],b+n(1-[e+\theta_{0}]))}=f(\theta_{1}|e,n),$
@@ A.2 Posterior Probability of $H_{1}$ for Two-arm Trial with Continuous outcome and known variance
$f(\theta|\bm{y}_{n}^{d})=\frac{1}{\sqrt{\frac{2\pi}{\frac{1}{b}+\frac{n}{\sigma^{2}}}}}\exp{\left\{-\frac{1}{2}\left[\frac{1}{b}+\frac{n}{\sigma^{2}}\right]\left(\theta-\frac{1}{\frac{1}{b}+\frac{n}{\sigma^{2}}}\left[\frac{a}{b}+\frac{n\bar{y}^{d}}{\sigma^{2}}\right]\right)^{2}\right\}}$
$=\frac{1}{\sqrt{\frac{2\pi}{\frac{1}{b}+\frac{n}{\sigma^{2}}}}}\exp{\left\{-\frac{1}{2}\left[\frac{1}{b}+\frac{n}{\sigma^{2}}\right]\left(\theta-\frac{1}{\frac{1}{b}+\frac{n}{\sigma^{2}}}\left[\frac{a}{b}+\frac{ne}{\sigma^{2}}\right]\right)^{2}\right\}}=f(\theta|e,n),$
@@ A.3 Posterior Probability of $H_{1}$ for Two-arm Trial with Binary and Count-data Outcomes
Therefore, according to Equation (3) in the manuscript, it follows that $\text{Pr}(H=H_{1}|\bm{y}_{n})=\text{Pr}(H=H_{1}|\bar{y}_{1},\bar{y}_{0},n)$.
@@ A.4 BESS Algorithm 2’
3:Set $n_{\text{min}}$ and $n_{\text{max}}$ as the smallest and largest candidate sample size.
@@
A.5 Additional Simulation Setup and Results in Section 4.2
Additional simulation details
@@
Additional results
For one-arm trials, we follow the steps in Section 4.2, with $e=0.4$, $\theta_{0}=0$, $\theta^{*}=0.3$, and $c=0.8$. The result is 15 patients for binary outcome, 71 for continuous outcome, and 25 for count outcome. Next, we obtain the coherence plot for the three outcomes through simulation with true parameter of $\theta_{1}=0.4$, which is shown in Figure A.1.
@@ A.6 Additional Simulation Setup and Results in Section 5
Table A.2 shows the estimated sample sizes via BESS and the oracle sample size with the standard SSE method for other outcome and trial types.
Trail Outcome Evidence Confidence type I power BESS Standard SSE
type type $e$ $c$ error rate $\alpha$ $1-\beta$ $n$ $n$
One-arm Binary 0.35 0.8 0.24 0.83 60 63
Continuous 0.4 0.8 0.20 0.50 71 71
Count 0.375 0.8 0.20 0.64 40 41
Two-arm Continuous 0.15 0.8 0.20 0.40 32 33
Count 0.3 0.7 0.27 0.88 51 51
@@
A.8 BESS Software Description
@@
    • Description: Compute confidence, i.e., the posterior probability of ($H_{1}$) for the three outcome types of one arm trial.
    • Output: Confidence, i.e., the posterior probability of $H_{1}$ given sample size $n$, evidence $e$, and clinically minimum effect size $\theta^{*}$.
@@
    • Description: BESS Algorithm 2 for binary outcomes in a two-arm trial.
    • Output: Sample size per arm for binary outcomes in a two-arm trial.
@@
    • Description: BESS Algorithm 1 for continuous outcomes with known variance in a two-arm trial.
    • Output: Sample size per arm for continuous outcomes with known variance in a two-arm trial.
@@
    • Description: Compute confidence, i.e., the posterior probability of $H_{1}$ for binary outcome in two-arm trial.
    • Output: Confidence, i.e., the posterior probability of $H_{1}$ given sample size $n$, the numbers of responders in treatment and control arms, and clinically minimum effect size $\theta^{*}$.
@@
    • Description: Compute confience, i.e., the posterior probability of $H_{1}$ for continuous outcome in a two-arm trial.
    • Output: Confidence, i.e., the posterior probability of $H_{1}$ given sample size $n$, difference in mean responses between treatment and control arms, and clinically minimum effect size $\theta^{*}$.
@@
    • Description: Compute confidence, i.e., the posterior probability of $H_{1}$ for count outcome in two-arm trial.
    • Output: Confidence, i.e., the posterior probability of $H_{1}$ given sample size $n$, event rates in treatment and control arms, and clinically minimum effect size $\theta^{*}$.